# Invariant Coordinate Selection and Fisher discriminant subspace beyond the case of two groups

### v1

@@ Appendix A.2.   The case of Dirac mixture in Section 3
The calculations are very similar to the ones in Appendix A.1 with the within-group matrix $\Gamma_{W}$ equals to the zero matrix, which removes some terms in the scatter expressions. In accordance with the previously established notation, it yields $\mathrm{COV}=[\beta_{ms}]$ where $\beta_{ms}=\sum_{j=1}^{k}\alpha_{j}t_{jm}^{\mbox{c}}t_{js}^{\mbox{c}}$ for $m,s\in\{1,\ldots,q\}$, and $\mathrm{COV}_{4}=\frac{1}{q+2}\times\mathbb{E}\Bigg{[}\Bigg{(}\sum_{i=1}^{q}\sum_{j=1}^{q}x^{\mbox{c}}_{i}x^{\mbox{c}}_{j}b_{ij}\Bigg{)}X^{\mbox{c}}(X^{\mbox{c}})^{\top}\Bigg{]}.$ The moments of a random variable following a Dirac distribution centered at $t^{\mbox{c}}$ is given by $m_{\nu}=(t^{{\mbox{c}}})^{\nu}$. All the moments can be expressed as $\mathbb{E}[x^{\mbox{c}}_{a}x^{\mbox{c}}_{b}x^{\mbox{c}}_{c}x^{\mbox{c}}_{d}]=\sum_{j=1}^{k}\alpha_{j}t^{\mbox{c}}_{ja}t^{\mbox{c}}_{jb}t^{\mbox{c}}_{jc}t^{\mbox{c}}_{jd}$ for $a,b,c,d\in\{1,\ldots,q\}$. It yields $\mathrm{COV}_{4}=[\psi_{ms}]$ where $\psi_{ms}=\frac{1}{q+2}\Bigg{[}\sum_{l=1}^{k}\alpha_{l}t^{\mbox{c}}_{lm}t^{\mbox{c}}_{ls}(t^{\mbox{c}}_{l})^{\top}\mathrm{COV}^{-1}(t^{\mbox{c}}_{l})\Bigg{]}$ for $m,s\in\{1,\ldots,q\}$.
$(\mathrm{COV}^{-1}\mathrm{COV}_{4})_{ms}=\frac{1}{q+2}\Bigg{[}\sum_{l=1}^{k}\alpha_{l}(t^{\mbox{c}}_{l})^{\top}\mathrm{COV}^{-1}(t^{\mbox{c}}_{l})\sum_{i=1}^{q}b_{mi}t^{\mbox{c}}_{li}t^{\mbox{c}}_{ls}\Bigg{]}$ for $m,s\in\{1,\ldots,q\}$.
Appendix B.   Details for the study of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ eigenvalues in Subsections 2.4, 3.2, 4.2 and additional figure
Subsection 2.4 and Sections 3 and 4 study the eigenvalues behavior of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$. Appendix B.1 provides details about the analysis in the case of a Gaussian mixture (Subsection 2.4) whereas Appendix B.2 explains the study in the case of a Dirac mixture (Sections 3 and 4).
Appendix B.1.   Gaussian mixture
In Subsection 2.4, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are analysed in the case of a mixture of a Gaussian distributions. Figure 1 displays the eigenvalues obtained with numerical computations which are based on the theoretical ones described in Appendix A.1. In order to comprehend the behavior of eigenvalues for any number of groups $k$ and any value of the dimension spanned by the group centers $q$, we will examine in particular the following cases: (i) $k=2$: $q=1$, (ii) $k=3$: $q=2$, (iii) $k=5$: $q=2$ and $q=4$. There are two cases with $k=5$ to illustrate the two subcases $q=k-1$ and $q<k-1$. To perform such numerical computations, we set to 6 the number of variables $p$, and we specify values for the group proportions and the group means. The latter are described in Table 1. We recall from Subsection 2.2 that:
$T=\begin{pmatrix}T_{u}&0\\ 0&0\\ \end{pmatrix},$
where $T=(t_{1},\ldots,t_{k})$ with $t_{j}$ a $p$-dimensional vector for $j\in\{1,\ldots,k\}$, and $T_{u}$ is an upper triangular matrix of dimension $k-1\geq 1$ such that the last $k-1-q\geq 0$ rows are zero. In Table 1, only the elements above the diagonal of $T_{u}$ are mentioned. The values are selected manually, ensuring that the matrix $T_{u}$ is full rank and that different configurations are represented. Some group means exhibit a particular structure. The first configuration of group means has a regular increment $\delta_{i}$ for each dimension, i.e. $t_{(j+1)i}=t_{ji}+\delta_{i}$ if $t_{ji}\neq 0$ for $j\in\{1,\ldots,k-1\}$, $i\in\{1,\ldots,p\}$ and $t_{ji}$ being the $i$-th element of $t_{j}$. The second configuration is equal to the first one multiplied by 10, i.e. $T_{2}=10\times T_{1}$ where $T_{1}$ ($T_{2}$) are the group mean matrices of the first (second) configuration. The third configuration is a diagonal matrix. The remaining group mean configurations were randomly selected to ensure a variety of configurations. The dimensions of the matrix containing the group means are determined by the number of groups. Consequently, the initial $(k-1)$ columns of Table 1 are selected, and a column of zeros is appended to the end. Only the first $q$-th rows are selected and zeros must be added up to the $p$-th row in order to fill the elements below the diagonal of $T_{u}$. This methodology yields the desired structure, which is consistent with the number of groups selected. For example, if $k=3$ and $p=3$, we obtain from the first row of Table 1 $T_{1,q=2}$ if $q=2$ and $T_{1,q=1}$ if $q=1$:
$T_{1,q=2}$ $=\begin{pmatrix}200&200&0\\ 0&400&0\\ 0&0&0\\ \end{pmatrix},$ $T_{1,q=1}$ $=\begin{pmatrix}200&400&0\\ 0&0&0\\ 0&0&0\\ \end{pmatrix}.$
For each value of $k$, we select mixture proportions, i.e., the values of $\alpha_{j}$ for $j\in\{1,\ldots,k\}$. The values have been selected to represent a range of scenarios and are listed in Table 3. For each number of groups, there is a scenario in which all groups have the same proportion. Then, the number of “low” and “large” groups is varied.
Next, for each mixture proportions scenario and each group centers configuration, the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ are calculated. The results are plotted on a graph with the group proportions on the $x$-axis and a boxplot per eigenvalue on the $y$-axis, representing the distribution of the eigenvalue across the group centers for a given scenario. This representation permits the comparison of eigenvalues across scenarios and across group centers for a given number of groups.
Appendix B.2.   Dirac Mixture
Section 3 analyses the eigenvalues of $V_{1}^{-1}V_{2}$ in the case of a mixture of a Dirac distributions where $p=q$. Figures 2, 10 and the additional Figure 10 show the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ obtained with numerical calculations which are based on the theoretical ones described in Appendix A.2. To understand the behavior of eigenvalues for any number of groups, we will examine the cases of a mixture of 2, 3, 5, and 10 groups in particular. For each of these cases, we select mixture proportions scenarios that are described in Table 4. For 2, 3 and 5 groups, we added more scenarios to the one from Table 3. For $k=10$, we created scenarios following the same logic. The centers of the groups are the same as those in Appendix B.1, described in Table 1. In Figure 2, $q=k-1$. In Figures 3 and 10, $q<k-1$. More precisely, we looked at: (i) $k=3$: $q=1$, (ii) $k=5$: $q=1$, $q=2$ and $q=3$, (iii) $k=10$: $q=1$, $q=3$, $q=5$ and $q=7$.
In Subsection 4.2, we still consider the case of a Dirac mixture where $p=q$. However, the subsection focuses on the case $q=k-1$ and does not address the impact of group centers. The group centers are then fixed to the configuration given by the first row of Table 1. Subsection 4.2 studies the idea of threshold values in the group proportions that result in an eigenvalue transitioning from less than one to more than one. Figure 7 illustrates these thresholds. Its configuration is very similar to the one of Figure 2: the represented number of groups $k$ are 2, 3, 5 and 10, and $p=q=k-1$. The proportions scenarios are described in Table 4. For each value of $k$, there is one scenario in which the proportion of the first group is at threshold level. The threshold is defined with the Setup 1 explained in Appendix C.2.
Table 1: Group centers used to compute the theoretical eigenvalues
(a) Groups 1 to 5
$t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$ $t_{5}$
1 [200] [400, 100] [600, 300, 200] [800, 500, 300, 300] [1000, 700, 400, 500, 200]
2 [2000] [4000, 1000] [6000, 3000, 2000] [8000, 5000, 3000, 3000] [10000, 7000, 4000, 5000,
2000]
3 [2] [0, 2] [0, 0, 6] [0, 0, 0, 7] [0, 0, 0, 0, 5]
4 [1] [5, 7] [9, 10, 2] [13, 13, 3, 4] [17, 16, 4, 8, 6]
5 [6] [-3, 7] [-10, 6, 4] [1, 2, 8, 6] [3, -3, 7, 9, 4]
6 [60] [-30, 70] [5, 15, 45] [75, 23, 54, 66] [59, 86, 38, 29, -32]
7 [18] [9, 14] [15, 22, 150] [65, 42, 32, 15] [22, 45, 12, 28, 41]
8 [-2] [11, 4] [0, 0, 10] [0, 3, 0, 13] [0, 0, 11, 34, 14]
9 [1200] [180, 910] [320, 112, 1000] [550, 321, 875, 200] [1000, 710, 593, 340, 900]
10 [36] [39, 66] [15, 88, 18] [48, 67, 36, 83] [22, 59, 48, -43, 36]
11 [300] [460, 180] [250, 80, 110] [410, 100, 230, 260] [230, 200, 160, 420, 318]
12 [12] [50, 27] [5, 13, 31] [25, 25, 31, 42] [49, 37, 21, 39, 45]
13 [16] [-30, 18] [1, 18, 50] [11, -18, 59, -36] [11, -18, 64, -12, 39]
14 [63] [-35, 12] [22, -12, 19] [55, -71, 22, 38] [54, -60, 31, 23, 45]
 15 [22] [9, 33] [61, 42, 41] [86, 42, 71, 30] [48, 67, 36, 83, 22]
16 [-2] [1, 2] [3, 4, 6] [8, 1, 5, 24] [3, -3, 7, 9, 5]
17 [3] [0, 6] [0, 0, 9] [0, 0, 0, 4] [0, 0, 0, 0, 5]
18 [400] [200, 550] [280, -450, 356] [312, 718, 462, 425] [410, 100, 230, 0, 300]
19 [1080] [2100, 1800] [1870, 3200, 2200] [1200, 2690, 3850, 4086] [4800, -1700, 600, 2300, 2200]
20 [54] [32, 41] [12, 33, 81] [43, 12, 79, 84] [11, 18, 59, 36, 23]
(b) Groups 6 and 7
$t_{6}$ $t_{7}$
1 [1200, 900, 500, 700, 300, 100] [1400, 1100, 600, 900, 400, 200, 100]
2 [12000, 9000, 5000, 7000, 3000, 1000] [14000, 11000, 6000, 9000, 4000, 2000, 1000]
3 [0, 0, 0, 0, 0, 8] [0, 0, 0, 0, 0, 0, -3]
4 [21, 19, 5, 12, 12, 10] [25, 22, 6, 16, 18, 20, 9]
5 [-4, 2, 8, -5, 12, 7] [4, 12, 5, -7, 23, 3, 8]
6 [77, 43, -55, 63, 0, 48] [54, 92, 19, 74, -23, 44, 32]
7 [12, 33, 16, 32, 46, 33] [17, 33, 21, 38, 57, 49, 67]
8 [0, 0, 8, 28, 17, 16] [0, 0, 17, 22, 31, 2, 16]
9 [370, 212, 2000, 560, 230, 410] [220, 230, 900, 120, 222, 714, 600]
10 [65, 45, 32, 56, 74, 43] [8, 43, 59, -23, 31, 32, 54]
11 [530, 210, 310, 450, 530, 310] [50, 100, 300, 536, 430, 320, 120]
12 [45, 27, 32, 65, 27, 33] [32, 37, 26, 34, 45, 43, 21]
13 [12, 37, 21, 22, 7, 12] [23, 37, 21, 34, 45, 32, 12]
14 [22, 37, 21, 38, 45, 69] [22, 33, 64, 62, 45, 23, 36]
15 [49, 37, 21, 39, 45, 13] [39, 45, 22, 45, 76, 23, 34]
16 [5, 2, 6, -4, 3, 8] [1, 15, 4, 5, 6, 7, -3]
17 [0, 0, 0, 0, 0, 6] [0, 0, 0, 0, 0, 0, 2]
18 [390, 0, 450, 360, -300, 380] [690, -212, 500, 560, -120, 410, 340]
19 [2100, 1000, -1000, 1389, 2300, 4500] [500, 1000, 0, 4362, 4300, 3200, 1200]
20 [49, 37, 21, 39, 7, 65] [45, 37, 32, 65, 28, 31, 66]
Table 2: Group centers used to compute the theoretical eigenvalues – Continued
(a) Groups 8 and 9
$t_{8}$ $t_{9}$
1 [1600, 1300, 700, 1100, 500, 300, 200, 100] [1800, 1500, 800, 1300, 600, 400, 300, 200, 100]
2 [16000, 13000, 7000, 11000, 5000, 3000, 2000, [18000, 15000, 8000, 13000, 6000, 4000, 3000, 2000,
1000] 1000]
3 [0, 0, 0, 0, 0, 0, 0, 16] [0, 0, 0, 0, 0, 0, 0, 0, 20]
4 [29, 25, 7, 20, 24, 30, 14, 5] [34, 27, 8, 24, 30, 40, 19, 10, 4]
5 [2, 32, 14, 3, -3, 0, 12, 10] [6, 21, 19, 8, 7, 11, 26, 20, 16]
6 [34, 56, 44, 22, 0, 38, 0, 32] [51, 32, 39, 30, 10, 78, 40, 28, 29]
7 [12, 20, 11, 31, 37, 9, 12, 30] [0, 5, 31, 16, 7, 11, 19, 20, 17]
8 [0, 0, 7, 17, 27, 13, 9, 13] [0, 0, 17, 32, 41, 5, 3, 19, 22]
9 [320, 530, 1100, 1200, 333, 743, 800, 900] [120, 130, 2200, 3200, 444, 743, 2100, 200, 300]
10 [18, 32, 14, 45, 89, 21, 43, 32] [20, 22, 44, 56, 61, 41, 61, 53, 33]
11 [110, 300, 500, 136, 230, 320, 0, 220] [210, 0, 650, 536, 560, 400, 30, 100, 300]
12 [12, 32, 25, 35, 51, 54, 31, 50] [32, 54, 76, 36, 22, 62, 66, 13, 56]
13 [37, 6, 54, 32, 0, 66, 10, 12] [19, 26, 32, 54, 0, 55, 34, 17, 31]
14 [-39, 32, 22, 81, 56, 38, 74, 12] [55, 34, 69, 65, 26, 38, 64, 42, 22]
15 [32, 38, 44, 52, 19, 42, 51, 21] [25, 8, 79, 63, 43, 72, 61, 51, 32]
16 [2, 20, 11, 3, 3, 9, 12, 10] [22, 2, 1, 4, 0, 5, 1, 8, 13]
17 [0, 0, 0, 0, 0, 0, 0, 8] [0, 0, 0, 0, 0, 0, 0, 0, 5]
18 [0, 89, 345, 280, 0, 270, 500, 220] [610, 0, 385, 680, 0 ,470, 76, 290, 300]
19 [0, 3100, 1400, 3800, 2300, -1100, 4200, 1800] [3900, 100, 4200,1900, 3700, 1100, 3100, -300, 2400]
20 [41, 32, 24, 74, 51, 23, 44, 65] [27, 6, 51, 42, 61, 56, 33, 55, 52]
Table 3: Scenarios analysed in Figure 1
k Scenarios
2 50-50, 20-80, 10-90
3 33-33-34, 10-40-50, 10-10-80
5 20-20-20-20-20, 10-10-20-20-40, 10-10-10-10-60
Fig. 10: Boxplots of the eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ with no within-group variability, where the group centers vary across 20 different configurations, with $k=10$ and $p=q<k-1$. The values of $q$ vary across the panels. The 12 group proportions scenarios ($\alpha_{j}$, $j\in\{1,\ldots,k\}$) vary on the $x$-axis.
Table 4: Scenarios analysed in Figures 2, 3, 7 and 10
k Scenarios for Figures 2, 3, 10 Scenarios for Figure 7
2 50-50, 40-60, 30-70, 20-80, 10-90 30-70, 21-79, 10-90
3 33-33-34, 20-30-50, 10-40-50, 10-30-60, 10-20-70, 10-10-80 33-33-34, 18-32-50, 10-40-50, 10-10-80
5 20-20-20-20-20, 10-20-20-20-30, 20-20-20-20-20, 14-20-20-20-26,
10-10-20-20-40, 10-10-10-30-40 10-10-20-20-40, 10-10-10-30-40
10 10-…-10, 5-5-5-10-10-10-10-10-15-20, 10-…-10, 8-10-10-10-10-10-10-10-10-12,
5-5-5-5-5-5-5-15-20-30 5-5-5-5-5-5-5-15-20-30
Appendix C.   Details for the study of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ eigenvalues when group proportions vary in the absence of within-group variability in Section 4
@@ Appendix B.2.   Dirac Mixture
Appendix C.1.   The case of three groups
@@
Subsection 4.1 focuses on the case where the number of groups $k$ is 3, and thus the number of variables $p$ is 2. The eigenvalues are represented in ternary diagrams in Figures 4 and 5. The construction of such figures is achieved through the generation of a grid comprising all possible combinations of the group proportions $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, with values between zero and one, such that their sum is equal to one. The step size is 0.001 (0.1%). This approach ensures exhaustive coverage of the proportion space and guarantees that every point on the ternary diagram represents a valid combination of $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$. The eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$, $\rho_{1}$ and $\rho_{2}$, are calculated for each combination of the grid. To perform this calculation, it is necessary to have values for the centers of the three groups. As these values will have no impact on the result, they may be assigned randomly. For the sake of simplicity, we choose the first row of Table 1.
@@
In Figure 4, it is of interest to highlight the case in which the eigenvalue is equal to one (white), as this is the case in which ICS does not “work”. Subsequently, two situations are distinguished: when the eigenvalue is smaller than one (blue), and when the eigenvalue is greater than one (red). Both blue and red parts follow a color gradient to illustrate the distance of the value from one. The logarithm of the eigenvalue was used in the coloring of the plot, because it allows for a more balanced distribution of colors. The logarithmic scale reduces the range difference between the blue and red parts, and since $log(1)=0$, it conveniently positions the transition color at a meaningful point. The scale of the gradient ranges from blue at the minimum of $log(\rho)$ to red at the maximum of $log(\rho)$, $log(\rho)=0$ being white. The gradient effectively illustrates the variations in the eigenvalue across the group proportions grid and highlights instances where ICS does not “work”. However, the color bar at the bottom of the plot represents the original eigenvalue, enhancing the plot’s interpretability.
@@
Both eigenvalues are shown qualitatively in Figure 5. Since the information of interest is whether the eigenvalues are greater or less than one (as explained in Subsection 4.2, three cases are distinguished: $\rho_{1}$ and $\rho_{2}>1$ (red), $\rho_{1}>1$ and $\rho_{2}<1$ (purple), and $\rho_{1}$ and $\rho_{2}<1$ (blue)).
Appendix C.2.   The case of $k$ groups

### v2

@@ Appendix A.2.   The case of Dirac mixture in Section 3
Upon the removal of the noise and the dimensions not associated with FDS, the Gaussian distributions are simplified to Dirac distributions with parameter $\mathbf{t}_{j}$, denoted by $\delta_{\mathbf{t}_{j}}$: $\mathbf{X}\sim\sum_{j=1}^{k}\alpha_{j}\delta_{\mathbf{t}_{j}},$ where $\mathbf{t}_{j}=\sum_{i=1}^{j}t_{ji}\mathbf{e}_{i}$, and $\mathbf{e}_{i}$ is a $q$-dimensional vector with one in the $i$-th coordinate and zero elsewhere for $j\in\{1,\ldots,k-1\}$, $\mathbf{t}_{k}$ is the zero vector, and the $\mathbf{t}_{j}$ are distinct for $j\in\{1,\ldots,k\}$. The calculations are very similar to the ones in Appendix A.1 with the within-group matrix $\boldsymbol{\Gamma}_{W}$ equals to the zero matrix, which removes some terms in the scatter expressions. In accordance with the previously established notation, it yields $\mathrm{COV}=[\beta_{ms}]$ where $\beta_{ms}=\sum_{j=1}^{k}\alpha_{j}t_{jm}^{\mbox{c}}t_{js}^{\mbox{c}}$ for $m,s\in\{1,\ldots,q\}$, and
$\mathrm{COV}_{4}=\frac{1}{q+2}\times\mathbb{E}\Bigg{[}\Bigg{(}\sum_{i=1}^{q}\sum_{j=1}^{q}x^{\mbox{c}}_{i}x^{\mbox{c}}_{j}b_{ij}\Bigg{)}\mathbf{X}^{\mbox{c}}(\mathbf{X}^{\mbox{c}})^{\top}\Bigg{]}.$
The moments of a random variable following a Dirac distribution centered at $t^{\mbox{c}}$ is given by $m_{\nu}=(t^{{\mbox{c}}})^{\nu}$. All the moments can be expressed as $\mathbb{E}[x^{\mbox{c}}_{a}x^{\mbox{c}}_{b}x^{\mbox{c}}_{c}x^{\mbox{c}}_{d}]=\sum_{j=1}^{k}\alpha_{j}t^{\mbox{c}}_{ja}t^{\mbox{c}}_{jb}t^{\mbox{c}}_{jc}t^{\mbox{c}}_{jd}$ for $a,b,c,d\in\{1,\ldots,q\}$. It yields $\mathrm{COV}_{4}=[\psi_{ms}]$ where
$\psi_{ms}=\frac{1}{q+2}\Bigg{[}\sum_{\ell=1}^{k}\alpha_{\ell}t^{\mbox{c}}_{\ell m}t^{\mbox{c}}_{\ell s}(\mathbf{t}^{\mbox{c}}_{\ell})^{\top}\mathrm{COV}^{-1}(\mathbf{t}^{\mbox{c}}_{\ell})\Bigg{]},m,s\in\{1,\ldots,q\},$
$(\mathrm{COV}^{-1}\mathrm{COV}_{4})_{ms}=\frac{1}{q+2}\Bigg{[}\sum_{\ell=1}^{k}\alpha_{\ell}(\mathbf{t}^{\mbox{c}}_{\ell})^{\top}\mathrm{COV}^{-1}(\mathbf{t}^{\mbox{c}}_{\ell})\sum_{i=1}^{q}b_{mi}t^{\mbox{c}}_{\ell i}t^{\mbox{c}}_{\ell s}\Bigg{]},m,s\in\{1,\ldots,q\}.$
Appendix B.   Details for the study of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$ eigenvalues when group proportions vary in the absence of within-group variability in Section 4
@@
Appendix B.1.   The case of three groups
@@
Subsection 4.1 focuses on the case where the number of groups $k$ is 3, and thus the number of variables $p$ is 2. The eigenvalues are represented in ternary diagrams in Fig. 4 and 5. The construction of such figures is achieved through the generation of a grid comprising all possible combinations of the group proportions $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, with values between zero and one, such that their sum is equal to one. The step size is 0.001 (0.1%). This approach ensures exhaustive coverage of the proportion space and guarantees that every point on the ternary diagram represents a valid combination of $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$. The eigenvalues of $\mathrm{COV}^{-1}\mathrm{COV}_{4}$, $\rho_{1}$ and $\rho_{2}$, are calculated for each combination of the grid. To perform this calculation, it is necessary to have values for the centers of the three groups. As these values will have no impact on the result, they may be assigned randomly. For the sake of simplicity, we choose the first row of Table S1.
@@
In Fig. 4, it is of interest to highlight the case in which the eigenvalue is equal to one (white), as this is the case in which ICS does not “work”. Subsequently, two situations are distinguished: when the eigenvalue is smaller than one (blue), and when the eigenvalue is greater than one (red). Both blue and red parts follow a color gradient to illustrate the distance of the value from one. The logarithm of the eigenvalue was used in the coloring of the plot, because it allows for a more balanced distribution of colors. The logarithmic scale reduces the range difference between the blue and red parts, and since $log(1)=0$, it conveniently positions the transition color at a meaningful point. The scale of the gradient ranges from blue at the minimum of $log(\rho)$ to red at the maximum of $log(\rho)$, $log(\rho)=0$ being white. The gradient effectively illustrates the variations in the eigenvalue across the group proportions grid and highlights instances where ICS does not “work”. However, the color bar at the bottom of the plot represents the original eigenvalue, enhancing the plot’s interpretability.
@@
Both eigenvalues are shown qualitatively in Fig. 5. Since the information of interest is whether the eigenvalues are greater or less than one (as explained in Subsection 4.1, three cases are distinguished: $\rho_{1}$ and $\rho_{2}>1$ (red), $\rho_{1}>1$ and $\rho_{2}<1$ (purple), and $\rho_{1}$ and $\rho_{2}<1$ (blue)).
Appendix B.2.   The case of $k$ groups